How much water must be added to 400L of mixture that is 80% alcohol to reduce it to a 60% mixture?
Answer by TimothyLamb(4379)

x = liters of 80% alcohol/water = liters of 20% water/alcohol
y = liters of pure water
z = liters of 60% alcohol/water = liters of 40% water/alcohol
---
x = 400
z = x + y
z = 400 + y
0.20x + 1.00y = 0.40z
0.20*400 + 1.00y = 0.40*(400 + y)
0.20*400 + y = 0.40*400 + 0.40y
y - 0.40y = 0.40*400 - 0.20*400
0.60y = 0.40*400 - 0.20*400
y = (0.40*400 - 0.20*400)/0.60
y = 133.333333333
---
answer:
y = liters of pure water = 133 1/3
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---